Ziggys yogurt blends reg. yogurt that is 3% fat with its no-fat yogurt to obtain low fat yogurt that is 1% fat. How many pounds of reg. yogurt and how many pound of no-fat yogurt should be mixed to obtain 60 pounds of low fat yogurt.
Answer by ptaylor(1894)

Let x=amount of no-fat yogurt needed
Then 60-x=amount of 3% fat yogurt needed
Now we know that the amount of pure fat in the no-fat yogurt(0) plus the amount of pure fat in the 3% yogurt (0.03)(60-x) has to equal the amount of pure fat in the final mixture (0.01)(60). So our equation to solve:
0.03(60-x)=0.01(60) get rid of parens
1.8-0.03x=0.6 subtract 1.8 from both sides
1.8-1.8-0.03x=0.6-1.8 collect like terms
-0.03x=-1.2 divide both sides by -0.03
x=40 lb---------------------------------------amount of no-fat yogurt needed
60-x=60-40=20 lb---------------------------------amount of reg yogurt needed
40*0+20*0.03=0.01*60
0.6=0.6

Hope this helps---ptaylor